How would you prepare 500ml of 28% w/v (weight/volume) KOH (potassium hydroxide) solution? From this stock, how would you prepare a 0.1M working solution?


1
Expert's answer
2022-04-19T03:11:59-0400

"m(KOH)=500\ ml\ *\ 0.28\ w/v=140\ g"

"n(KOH)=\dfrac{Mass}{Molar\ mass}=\dfrac{140\ g}{56\ g/mol}=2.5\ mol"

"c(KOH)=\dfrac{Moles}{Volume}=\dfrac{2.5\ mol}{0.5\ L}=5M"


To prepare a 0.1M solution we need to dilute "\dfrac{5M}{0.1M}=50"-fold with water. If 500 mL of working solution is needed, then "\dfrac{500\ mL}{50}=10\ mL" of the stock solution should be diluted with water to 500 mL.